# Meta-Homogenization for Knitwear Simulation

### v2

@@ 8. Experimental Results
We implemented our meta-homogenization framework on a desktop computer with an Intel i7-12700 CPU and an NVIDIA RTX 3090 GPU. We used Spectra library for computing the eigendecomposition of the mesh Laplacian, and $\bm{K}^{ii}_{d}$ for each domain, and Eigen (Guennebaud et al., 2010) as the primary interface of linear algebra computations. The domain-decomposed PD procedure consists of two steps. The first step solves the global matrix in the subspace, and we employed the PARDISO solver (Schenk et al., 2001) shipped with Intel MKL (Wang et al., 2014). Fullspace A-Jacobi iterations (Lan et al., 2022) are followed to make sure the global step system is well solved. This step was implemented on the GPU with CUDA. We used Chebyshev weight to improve the convergence of A-Jacobi method (Wang, 2015). The local step is parallelized on the GPU at each element with CUDA. Adjoint gradient descent and adjoint Gauss-Newton are sensitive to the residual of Eq. 15 i.e., Eq. 15 must be strictly satisfied. Therefore, during the elasticity fitting, we applied a few (three to five) Newton iterations after the PD solve to keep the residual smaller than $51E-$ 5. We generate full-scale YLS results by simulating each yarn as the Cosserat rod model (Spillmann and Teschner, 2007). The detailed statistics of experiments are reported in Tab. 1.
@@
8.1. Comparison with HYLC
Homogenized yarn-level cloth (HYLC) (Sperl et al., 2020) provides an excellent paradigm by homogenizing the “yarn material” to the mid-surface of the fabric, represented as a triangle mesh. Our meta-homogenization method expands the dimensionality to a 3D volume with more DOFs and optimizes spatially varying material properties instead of assuming a uniform material as in HYLC. Our first experiment involves three side-by-side comparisons to thoroughly understand the differences between our method and HYLC  [^1] We used the implementation of HYLC at https://git.ista.ac.at/gsperl/HYLC.
As shown in Fig. 7, both HYLC and our method yield visually similar and plausible results for a knitted patch with a 1$\times$1 rib pattern compared to YLS as the ground truth. The 1$\times$1 rib pattern is highly stretchable due to the alternating knit and purl columns, allowing the fabric to expand and contract significantly across its width. However, HYLC, utilizing a single material across the entire mesh, causes the patch to narrow in the perpendicular directions when stretched, exhibiting a rubber-like behavior (see the rightmost column in Fig. 7). The inconsistency between our method and HYLC becomes more apparent for the stockinette patch, which is less stretchy than the rib but tends to curl at the edges. While both sheet-based (HYLC) and volume-based (ours) materials display curly edges under stretch, HYLC is less expressive in capturing subtle deformation variations across the fabric due to the lack of material heterogeneity. In contrast, our method better replicates this phenomenon. We further test by draping a 1$\times$1 rib patch over a sphere, demonstrating a similar difference to that shown in Fig. 8. Both HYLC and our method produce reasonably good global deformation. However, our method, employing a spatially varying material, better captures the inhomogeneous deformation at the contacting region between the fabric and the sphere. In this set of experiments, only two or three $\bm{x}^{Y}_{i}$ are used for elasticity fitting, and the draping simulation results are not observed during the elasticity fitting.
@@ 8.2. More patterns
We further compare our results with yarn-level simulation (YLS) using three additional knit patterns: basketweave, montague, and flame ribbing. Particularly, flame ribbing patterns exhibit anisotropic spatial variation with non-periodic yarn structures, making it challenging to fit a homogeneous material model using HYLC. In contrast, our meta-homogenization does not rely on any assumptions about the underlying arrangement of yarn structures. As shown in Fig. 10, our method produces high-fidelity knit stretching results that are nearly identical to YLS. As visualized on the right, the fitted material distributions align well with the underlying yarn structure and capture location-dependent deformations as expected.
@@ 8.3. Convergence
The key to a successful meta-homogenization is to solve the inverse problem using high-order optimization techniques i.e., the adjoint Gauss-Newton method discussed in Sec. 6. We note that most existing gradient-based methods, either using adjoint method or using AD (automatic differentiation) fail to converge in our case. We plot the convergence curves using adjoint Gauss-Newton and gradient descent for a representative elasticity fitting instance and show the result in Fig. 12. We also visualize the deformed mesh that satisfies Eq. 15 at three different iterations. A consistent observation is that first-order methods are never going to work. They frequently get trapped at local minima. A more severe issue is the step size of the gradient-based method is not stable. Performing line search is expensive in elasticity fitting – any proposed material update $\Delta\bm{\gamma}$ can only be checked when Eq. 15 is satisfied, and therefore a forward simulation procedure is invoked. This makes the first-order method prohibitive in practice. For instance, it could take over one week if we choose to use the first-order adjoint method or differentiable simulation to optimize the rib pattern (Fig. 7). On the other hand, Gauss-Newton finishes the training in about three hours.
@@
Figure 15. Puffer ball on the knit.  We can separate the processing of energy constraint (on the mesh) and collision constraint (at the yarn level) so that the homogenized model interacts like a yarn-level model. In this example, we show a puffer ball falls on a stockinette fabric. The hairs on the puffer ball pass through the gaps on the knit. They are two-way coupled under collision constraints and can be conveniently processed with our simulation.
Figure 16. Short- and long-sleeve Yangge.  In addition to Fig. 1, we show two more examples of full garment animations of the Yangge dance. On the left, the character wears a knitted short-sleeve sweater. It has a different pattern (stockinette pattern) than the one shown in Fig. 1, which yields different garment dynamics. On the right, the character wears a long-sleeve sweater in a stockinette pattern. The material distributions are visualized on the right.
@@ 8.4. More Results
We test our method in several complex scenes. Figs. 13 and 14 report two challenging simulation scenarios involving intensive twisting and bending. In Fig. 13, we rotate two fabric strips (basketweave) for about $1,920$ degrees. They are tightly intertwined to form a knot. In this case, meta-homogenization yields a stronger volume-preserving penalty than examples shown in Fig. 10. We would like to mention that while the homogenized simulation uses the tetrahedron mesh for energy integration and elasticity constraint projection, we can still process collision at yarn segments if needed (e.g., see Fig. 15). As most yarn-yarn contacts are inelastic, we follow the collision projection strategy as in vanilla PD (Bouaziz et al., 2014) for each edge-edge collision. CCD is used to make sure we do not miss inter-penetrating edge pairs. However, we do not use the interior-point method as in (Li et al., 2021), and we only apply the line search filtering at the end of a time step. In other words, our mesh-level simulation does not have guarantees and may generate inter-penetration eventually. Nevertheless, we find our homogenized animations are visually plausible. Fig. 14 showcases an animation of twisting a wider fabric. Similar to Fig. 13, our method faithfully captures the twisting behaviors of such a complicated yarn structure using our volume-preserving energy. The YLS result can be found in the supplementary video. There are $91$K elements in Fig. 13 and $204$K elements in Fig. 14. The simulation uses $335$ ms for one time step on average in Fig. 13 and $672$ ms in Fig. 14.
@@ 9. Conclusion & limitation
This paper explores a different perspective on improving the simulation of knitwear with complex knitting patterns. We name our method meta-homogenization because it enables a volumetric and spatially heterogeneous material synthesis. The advantages of this strategy are multifold. Volume materials implicitly handle the bending and twisting and reduce the nonlinearity of the material property. It allows a faster simulation algorithm and, in turn, benefits the fitting efficiency. Assigning each volume element an independent set of material parameters effectively enhances the versatility of simulation so that our method is able to capture subtle and visually pleasing local deformations of complex knittings. With a domain-decomposed PD solver, our method is orders of magnitude faster than full-scale YLS.
While meta-homogenization shows make some non-trivial improvements over existing methods, it still has drawbacks and limitations. We leverage shape fitting to extract the acceleration of YLS sequences and to isolate the inertia effects during the material learning. This works well in general. For fast-moving scenes however, the estimated inertia and lumped mass may still differ from the reference. Yarn-level cloth models are often highly dampened. This is because the friction and contacts among yarn threads dissipate inertia energy quickly. It is challenging to calibrate a homogenized garment with full-scale YLS using commonly seen macroscopic damping models e.g., Rayleigh damping. A dedicated data-driven damping model may be a better choice to overcome this difficulty. Being a data-driven method, our method could generate different results of the same knitwear given different YLS inputs. It is possible to learn a more sophisticated strain-stress model as in (Sperl et al., 2020) at each element. Doing so requires significant computations, and the resulting material becomes strongly nonlinear (again). In other words, it remains an open question to find the right compromise between material nonlinearity, the level of visual realism, and computational efficiency.

### v3

@@ 8. Experimental Results
We implemented our volumetric homogenization framework on a desktop computer with an Intel i7-12700 CPU and an NVIDIA RTX 3090 GPU. We used Spectra library for computing the eigendecomposition of the mesh Laplacian, and $\bm{K}^{ii}_{d}$ for each domain, and Eigen (Guennebaud et al., 2010) as the primary interface of linear algebra computations. The domain-decomposed PD procedure consists of two steps. The first step solves the global matrix in the subspace, and we employed the PARDISO solver (Schenk et al., 2001) shipped with Intel MKL (Wang et al., 2014). Fullspace A-Jacobi iterations (Lan et al., 2022b) are followed to make sure the global step system is well solved. This step was implemented on the GPU with CUDA. We used Chebyshev weight to improve the convergence of A-Jacobi method (Wang, 2015). The local step is parallelized on the GPU at each element with CUDA. Adjoint gradient descent and adjoint Gauss-Newton are sensitive to the residual of Eq. 14 i.e., Eq. 14 must be strictly satisfied. Therefore, during the elasticity fitting, we applied a few (three to five) Newton iterations after the PD solve to keep the residual smaller than $51E-$ 5. We generate full-scale YLS results by simulating each yarn as the Cosserat rod model (Spillmann and Teschner, 2007). The detailed statistics of experiments are reported in Tab. 1.
8.1. Comparison with HYLC
Homogenized yarn-level cloth (HYLC) (Sperl et al., 2020) provides an excellent paradigm by homogenizing the “yarn material” to the mid-surface of the fabric, represented as a triangle mesh. Our volumetric homogenization method expands the dimensionality to a 3D volume with more DOFs and optimizes spatially varying material properties instead of assuming a uniform material as in HYLC. Our first experiment involves three side-by-side comparisons to thoroughly understand the differences between our method and HYLC  [^1] We used the implementation of HYLC at https://git.ista.ac.at/gsperl/HYLC. In HYLC, the Discrete Elastic Rod energy is replaced by Cosserat Rod energy for our comparison. The visualization of both our method and HYLC is achieved by embedding the yarn geometry into the mesh.
As shown in Fig. 7, both HYLC and our method yield visually similar and plausible results for a knitted patch with a 1$\times$1 rib pattern compared to YLS as the ground truth. The 1$\times$1 rib pattern is highly stretchable due to the alternating knit and purl columns, allowing the fabric to expand and contract significantly across its width. However, HYLC, utilizing a single material across the entire mesh, causes the patch to narrow in the perpendicular directions when stretched, exhibiting a rubber-like behavior (see the rightmost column in Fig. 7). The inconsistency between our method and HYLC becomes more apparent for the stockinette patch, which is less stretchy than the rib but tends to curl at the edges. While both sheet-based (HYLC) and volume-based (ours) materials display curly edges under stretch, HYLC is less expressive in capturing subtle deformation variations across the fabric due to the lack of material heterogeneity. In contrast, our method better replicates this phenomenon. We further test by draping a pre-stretched 1$\times$1 rib patch over a sphere, demonstrating a similar difference to that shown in Fig. 8. Both HYLC and our method produce reasonable global deformation. However, our method, with a spatially varying material, better captures the inhomogeneous deformation at the contacting region. In this set of experiments, only two or three $\bm{x}^{Y}_{i}$ are used for elasticity fitting, and the draping simulation results are not observed during the elasticity fitting.
@@ 8.1. Comparison with HYLC
[…]
@@ 8.2. More patterns
We further compare our results with yarn-level simulation (YLS) using three additional knit patterns: basketweave, montague, and flame ribbing. Particularly, flame ribbing patterns exhibit anisotropic spatial variation with non-periodic yarn structures, making it challenging to fit a homogeneous material model using HYLC. In contrast, our volumetric homogenization does not rely on any assumptions about the underlying arrangement of yarn structures. As shown in Fig. 10, our method produces high-fidelity knit stretching results that are nearly identical to YLS. As visualized on the right, the fitted material distributions align well with the underlying yarn structure and capture location-dependent deformations as expected.
@@ 8.3. Convergence
The key to a successful volumetric homogenization is to solve the inverse problem using high-order optimization techniques i.e., the adjoint Gauss-Newton method discussed in Sec. 6. We note that most existing gradient-based methods, either using adjoint method or using AD (automatic differentiation) fail to converge in our case. We plot the convergence curves using adjoint Gauss-Newton and gradient descent for a representative elasticity fitting instance and show the result in Fig. 12. We also visualize the deformed mesh that satisfies Eq. 14 at three different iterations. A consistent observation is that first-order methods are never going to work. They frequently get trapped at local minima. A more severe issue is the step size of the gradient-based method is not stable. Performing line search is expensive in elasticity fitting – any proposed material update $\Delta\bm{\gamma}$ can only be checked when Eq. 14 is satisfied, and therefore a forward simulation procedure is invoked. This makes the first-order method prohibitive in practice. For instance, it could take over one week if we choose to use the first-order adjoint method or differentiable simulation to optimize the rib pattern (Fig. 7). On the other hand, Gauss-Newton finishes the training in about three hours.
@@
[…]
Figure 15. Puffer ball on the knit.  We separate the processing of energy constraint (on the mesh) and collision constraint (at the yarn level) so that the homogenized model interacts as a yarn-level model. In this example, a puffer ball falls on a stockinette fabric. The hairs on the puffer ball pass through the gaps on the knit. They are two-way coupled under collision constraints and can be conveniently processed with our simulation.
[…]
Figure 16. Short- and long-sleeve Yangge.  In addition to Fig. 1, we show two more examples of full garment animations. On the left, the character wears a knitted short-sleeve sweater. It has a different pattern (stockinette pattern) than the one shown in Fig. 1, which yields different garment dynamics. On the right, the character wears a long-sleeve sweater in a stockinette pattern. The material distributions are visualized on the right.
[…]
8.4. Comparision at Different Resolution
Figure 17. Comparison at different resolution.  We compare our method’s fitting capabilities across varying mesh resolutions by presenting an instance of a basketweave fabric draped over a sphere. Each panel displays the fabric fitting result at a different level of resolution. The smallest thumbnail inset in the bottom left panel provides the ground truth. The middle and the leftmost panels, where each cell’s size is similar to or smaller than the size of each loop structure of threads, show fitting results that are essentially consistent with the ground truth. However, the rightmost panel, with a much coaser resolution, displays anisotropic bending effects macroscopically but significantly loses sharp bending details.
[…]
The resolution of our volumetric material significantly influences the fitting capability of our method. We compare our method’s performance on basketweave material at various resolutions as shown in Fig. 17. The center of the basketweave patch was fixed with a sphere. While the boundary condition exhibits rotational symmetry, the basketweave demonstrates distinct bending behaviors between the course and wale directions. When the mesh resolution is close to or higher than that of the knit pattern, our approach produces results nearly identical to the ground truth. When the mesh resolution is significantly lower than that of the knit pattern, we still observe anisotropic bending behavior, but the distinction between the bending behaviors in the course and wale directions is not as sharp as in the ground truth.
@@ 8.5. More Results
We test our method in several complex scenes. Figs. 13 and 14 report two challenging simulation scenarios involving intensive twisting and bending. In Fig. 13, we rotate two fabric strips (basketweave) for about $1,920$ degrees. They are tightly intertwined to form a knot. In this case, volumetric homogenization yields a stronger volume-preserving penalty than examples shown in Fig. 10. We would like to mention that while the homogenized simulation uses the tetrahedron mesh for energy integration and elasticity constraint projection, we can still process collision at yarn segments if needed (e.g., see Fig. 15). As most yarn-yarn contacts are inelastic, we follow the collision projection strategy as in vanilla PD (Bouaziz et al., 2014) for each edge-edge collision. The collision impulse is then interpolated to the corresponding tetrahedron (Lan et al., 2022a). For collision detection, to enhance performance, we adopt the DCD (Discrete collision detection) strategy, noting that edge-edge collision pairs that exist in the same cell or adjacent cells can be excluded in advance. Of course, if we aim for strict non-penetration, we can adopt CCD (Continuous collision detection) with the interior point method as in  (Li et al., 2021). Although in our experiments we used the DCD collision detection method, which does not have guarantees and may eventually result in inter-penetration, we find that our homogenized animations remain visually plausible. Fig. 14 showcases an animation of twisting a wider fabric. Similar to Fig. 13, our method faithfully captures the twisting behaviors of such a complicated yarn structure using our volume-preserving energy. The YLS result can be found in the supplementary video. There are $91$K elements in Fig. 13 and $204$K elements in Fig. 14. The simulation uses $335$ ms for one time step on average in Fig. 13 and $672$ ms in Fig. 14.
@@
8.6. Generalizability
Figure 18. Mixed patterns (draping).  We drape the mixed pattern on a cylinder, reusing the material of the mixed pattern trained in Fig. 11 Our method exhibits almost the same behavior as YLS.
[…]
Figure 19. More poses.  We follow the data generation protocol outlined in HYLC to sample five types of deformation, including their symmetric counterparts. The first two types are bending along the central axis and diagonal axis. The third deformation is a fan-like expansion. For the fourth type, the fabric is subject to a twist that forms a helix, we include the multi-folded deformation featuring several folds along the material.
[…]
Figure 20. Dropping to ground.  We utilize the training data generated from Fig. 19 and compare the dropping behavior with different training sets. Initially, only the first two bending types are used as our training data. In the subsequent panel, the fan-like expansion and twist types are also included. All deformation types in Fig. 19 are added in the third panel. The corresponding material distributions are shown in the top-right corner of each panel. The YLS is displayed in the rightmost column.
[…]
Most of our experiment’s training data is directly sampled from the animation sequence. However, in Fig. 9, the draping test of the pre-stretched patch uses training data based on lateral stretching deformation. The final result shows a similar draping behavior to YLS, demonstrating the generalizability of our method to some extent. Here, we provide additional experiments to further explore the generalizability of our method. Fig. 18 demonstrates the draping of mixed patterns on a cylindrical form, using the material of the mixed pattern trained as shown in Fig. 11, the homogenized simulation exhibits almost the same behavior. Besides sample training data from the animation sequence, we can also follow the data generation protocol as in HYLC to sample possible deformations. In Fig. 19, we generate five types of deformation poses: Panel 1 shows a simple upward bend along the central axis, creating a smooth concave shape; Panel 2 presents a diagonal bend, where the material curves along the diagonal axis; Panel 3 illustrates a fan-like stretch configuration; Panel 4 depicts a twisting deformation that forms a helicoidal structure; and Panel 5 features a more complex, multi-folded bend with several folds along the material. Each of these poses, along with their symmetric counterparts, progressively composes the training set used to train the material distribution, as shown in Fig. 20. As the training data set increases, the behavior becomes closer to that of YLS. It is important to note that the training dataset does not include any data from the YLS animation sequence shown in Fig. 20.
@@ 9. Conclusion & limitation
This paper explores a different perspective on improving the simulation of knitwear with complex knitting patterns. We name our method volumetric homogenization because it enables a volumetric and spatially heterogeneous material synthesis. The advantages of this strategy are multifold. Volume materials implicitly handle the bending and twisting and reduce the nonlinearity of the material property. It allows a faster simulation algorithm and, in turn, benefits the fitting efficiency. Assigning each volume element an independent set of material parameters effectively enhances the versatility of simulation so that our method is able to capture subtle and visually pleasing local deformations of complex knittings. With a domain-decomposed PD solver, our method is orders of magnitude faster than full-scale YLS.
While volumetric homogenization shows make some non-trivial improvements over existing methods, it still has drawbacks and limitations. We leverage shape fitting to extract the acceleration of YLS sequences and to isolate the inertia effects during the material learning. For fast-moving scenes however, the estimated inertia and lumped mass may still differ from the reference. Yarn-level cloth models are often highly dampened. This is because the friction and contacts among yarn threads dissipate inertia energy quickly. It is challenging to calibrate a homogenized garment with full-scale YLS using commonly seen macroscopic damping models e.g., Rayleigh damping. Being a data-driven method, our method could generate different results of the same knitwear given different YLS inputs. It is possible to learn a more sophisticated strain-stress model as in (Sperl et al., 2020) at each element. Doing so requires significant computations, and the resulting material becomes strongly nonlinear (again). In other words, it remains an open question to find the right compromise between material nonlinearity, the level of visual realism, and computational efficiency.